# Impact of electronic correlations on the superconductivity of high-pressure CeH9

### v1

@@ Conclusion
In summary, we have used a combination of DFT and DFMT to study electrons and phonons of CeH9 under high pressure, demonstrating that strong electronic correlations arising from cerium $4f$-states play an important role in this system. Many-body effects increase electron masses, produce a larger spectrum at the Fermi level, and soften hydrogen vibrational modes. Combining DFT+DMFT with the Migdal-Eliashberg theory of superconductivity, we were able to investigate the role of electronic correlations for the critical temperatures of phonon-mediated superconductivity in CeH9. Our findings indicate that the dominant correction to the critical temperature comes from the many-body renormalization of the electronic structure. However, we find that the renormalization of phonons is significant and cannot be neglected. The many-body-driven change in the critical superconducting temperature is 50 K. Future extensions of this work should address the impact of correlations on the el-ph coupling. In addition, previous works indicate that non-local exchange and correlation effects can increase the el-ph coupling [64, 65]. Overall, our findings shed new light on phonon-mediated high-temperature superconductivity in rare-earth superhydrides.
@@ Methods
We have performed DFT + DMFT calculations at inverse temperature $\beta=20$ eV-1 (i.e. $T\simeq 580$ K) with the embedded DMFT functional package Edmft [32] integrated with the all-electron DFT code Wien2k [66]. The DMFT impurity problem is solved using a continuous-time quantum Monte Carlo algorithm [37, 61]. We use the nominal double counting scheme [37], with an occupation fixed to 2, as the DFT-calculated valency of Ce under pressure is found to be 1.48 (greater than the nominal 1 electron) indicating an increased $f$-electron occupancy compared to the idealized trivalent state. A brief summary of Haule et al.’s implementation [32] is given in SI Sec. S5. We analytically continue the Matsubara self-energy to real frequencies using the maximum entropy method [67]. The quasiparticle weight was extracted from both Matsubara and real-frequency data and was found to be consistent. We have cross-checked the Wien2k all-electron results with the pseudopotential-based DFT packages CASTEP [68, 69] and Quantum ESPRESSO [70, 71] that were used previously in Ref. [59], using the same ultrasoft pseudopotentials that CASTEP generates on the fly for both codes. The valence electron configurations of $5s^{2}5p^{6}4f^{1}5d^{1}6s^{2}$ (with $4f$-electrons explicitly included) and $1s$ are adopted for the Ce and H atoms, respectively. Figure S9 in SI Sec. S6 shows a favorable comparison. Convergence tests at the DFT level yielded an energy cut-off of 800 eV for CASTEP and for Quantum ESPRESSO. For Wien2k, convergence is achieved when the product of the smallest atomic‐sphere radius and the maximum reciprocal‐space vector in the plane‐wave expansion reaches 7. A $\bm{\Gamma}$-centered $\mathbf{k}$-point grid of size $12\times 12\times 6$ is used for all three packages. The exchange-correlation functional is approximated using PBE [72]. As was done in previous works [12, 9, 47, 13, 59], we omit spin-orbit coupling for technical reasons. In the DFT + DMFT calculations, we use an on-site Coulomb interaction (Hubbard $U$) of $U=6.0$ eV and a Hund’s coupling $J=0.7$ eV for the cerium $f$ orbitals. This choice is comparable to previous values that were successfully used for elemental cerium [35, 36, 37, 38], CeH9 [59], and other Ce-compounds [44]. Note that we neglect any potential impact of pressure on Hubbard $U$ [41, 73]. We also omit any additional renormalizations of the DMFT electronic structure that arise from the electron-phonon coupling [74, 75].
@@
For the lattice dynamics calculations, we use finite difference with an in-house code which offers a versatile interface with DFT and DFT+DMFT [51], to calculate the phonon spectrum of CeH9. Considering that the computational cost of DFT+DMFT exhibits a particularly unfavorable scaling with the system size, we employ the non-diagonal supercell technique [76] to sample a $2\times 2\times 2$ $\mathbf{q}$-point grid in the Brillouin zone and then interpolate the result to obtain dynamical matrices on a $\mathbf{q}$-point grid of $6\times 6\times 3$. Specifically, we use four different supercells constructed with the following supercell matrices: $\big{(}\begin{smallmatrix}1&0&0\\ 0&1&0\\ 0&0&1\\ \end{smallmatrix}\big{)}$, $\big{(}\begin{smallmatrix}1&0&0\\ 0&1&0\\ 0&0&2\\ \end{smallmatrix}\big{)}$, $\big{(}\begin{smallmatrix}1&0&0\\ 0&2&0\\ 0&0&1\\ \end{smallmatrix}\big{)}$, and $\big{(}\begin{smallmatrix}1&0&0\\ 0&1&-1\\ 1&2&0\\ \end{smallmatrix}\big{)}$, the first one provides access to $\bm{\Gamma}-\mathbf{q}$-point, then the second one provides access to the $(0,0,1/2)$ $\mathbf{q}$-point, the third one to the $(0,1/2,0)$ $\mathbf{q}$-point, and the forth one to the $(1/2,1/2,1/2)$ $\mathbf{q}$-point. All other $\mathbf{q}$-points on the $2\times 2\times 2$ grid are related to these four by symmetry and therefore explicit calculations are not necessary.
@@
To evaluate the superconducting critical temperature $T_{\mathrm{c}}$, we used EPW [77, 78] to interpolate the el-ph coupling matrix calculated with the density functional perturbation theory as implemented in Quantum ESPRESSO from a $12\times 12\times 6$ $\mathbf{k}$-point grid and a $6\times 6\times 3$ $\mathbf{q}$-point grid to a $48\times 48\times 24$ $\mathbf{k}$-point grid and a $24\times 24\times 12$ $\mathbf{q}$-point grid. In particular, for $T_{\mathrm{c}}$ calculations at the DFT+DMFT level, the electronic and phononic degrees of freedom obtained from standard DFT are replaced with the counterparts obtained from DFT+DMFT. We used a standard value for the Morel-Anderson pseudopotential $\mu_{\mathrm{c}}=0.13$ which accounts for Coulomb repulsion of the system [56] in all our calculations. We present the Eliashberg theory of superconductivity as implemented in EPW in SI Sec. S7.
@@
The data that support the findings of this study are available in the Supplementary Information and on the Materials Cloud Archive.
@@
The Wien2k code used in this study is a commercial software, available from http://susi.theochem.tuwien.ac.at. The Quantum ESPRESSO code used in this research is open source: https://www.quantum-espresso.org. The CASTEP code used in this study is freely available at https://www.castep.org for academic research. The EPW code used in this research is open source: https://epw-code.org. The ana_cont code used in this study is open source: https://github.com/josefkaufmann/ana_cont.
@@
S.C. and Y.W. contributed equally to this work. S.C. performed the DMFT spectral function and superconducting calculations, while Y.W. performed the DMFT phonon calculations. S.P. and J.M.T provided technical help with the Wien2k and EPW software. S.P., B.M. and J.M.T supervised the work. All authors reviewed and approved the final version of the manuscript.

### v2

@@ Discussion
In summary, we have used a combination of DFT and DFMT to study electrons and phonons of CeH9 under high pressure, demonstrating that strong electronic correlations arising from cerium $4f$-states play an important role in this system. Many-body effects increase electron masses, produce a larger spectrum at the Fermi level, and soften hydrogen vibrational modes. Combining DFT+DMFT with the Migdal-Eliashberg theory of superconductivity, we were able to investigate the role of electronic correlations for the critical temperatures of phonon-mediated superconductivity in CeH9. Our findings indicate that the dominant correction to the critical temperature comes from the many-body renormalization of the electronic structure. However, we find that the renormalization of phonons is significant and cannot be neglected. The many-body-driven increase in the critical superconducting temperature is 50 K. Future extensions of this work should address the impact of many-body correlations on the perturbed potential, given that exchange and correlation effects beyond semi-local DFT can either increase [64, 65, 66, 67, 68] or decrease [69, 70] the values of the el-ph coupling matrix elements.
Overall, the renormalized Migdal-Eliashberg approach developed here is not specific to CeH9 but can be applied to other compounds where el-ph interactions coexist with electronic correlations. Promising targets include transition-metal hydrides with partially filled $d$-shells, such as X2MH6 (where X are alkali, alkaline-earth, or post-transition metals and M are 3d-, 4d-, and 5d-transition metals) which have recently attracted attention for their potential high-temperature superconductivity at ambient pressure [71]. In addition, while our approach focuses on phonon-mediated superconductivity, it could serve as a reference point to assess the relative importance of el-ph coupling effects in materials where other mechanisms such as spin-fluctuations [72, 73] dominate. Future developments incorporating electronic self-energy effects beyond the Migdal-Eliashberg theory may help clarify how phonons influence unconventional superconductivity in these systems.
@@ Methods
We have performed DFT + DMFT calculations at inverse temperature $\beta=20$ eV-1 (i.e. $T\simeq 580$ K) with the embedded DMFT functional package eDMFT [32] integrated with the all-electron DFT code WIEN2k [74] for both electrons and phonons in CeH9. A single calculation at a lower temperature with $\beta=100$ eV-1 (i.e. $T\simeq 116$ K) has also been performed to examine possible effects of temperature onto the electronic self-energy. The resulting self-energy was found to fall on top of that obtained at $\beta=20\penalty 10000\ \mathrm{eV}^{-1}$ (see Fig. S6a). The DMFT impurity problem is solved using a continuous-time quantum Monte Carlo algorithm [37, 62]. We use the nominal double counting scheme [37], with an occupation fixed to 2, as the DFT-calculated valency of Ce under pressure is found to be 1.42 (greater than the nominal 1 electron) indicating an increased $f$-electron occupancy compared to the idealized trivalent state. The impact of the choice of the nominal occupation values on the self-energy has been examined in Sec. S5 of the SI, where the quasiparticle renormalization factor is found to be independent when the nominal occupation varies from 1 to 2. The corresponding DMFT occupation is increasing from 1.11 to 1.58 for the same range. The standard nominal occupation for (unpressurized) Ce-compounds, $n_{f}^{0}=1$, corresponds to a Ce3+ configuration for pressurized CeH9, which we find implausible given the DFT valency for CeH9. We also note that the magnetic susceptibility of CeH9 exhibits no long-term memory, see SI Sec. S6, indicating the absence of fluctuating local moments in CeH9. Details of Haule et al.’s implementation can be found in Ref. [32]. We analytically continue the Matsubara self-energy to real frequencies using the maximum entropy method [75]. The quasiparticle weight was extracted from both Matsubara and real-frequency data and was found to be consistent. We have cross-checked the Wien2k all-electron results with the pseudopotential-based DFT packages CASTEP [76, 77] and Quantum ESPRESSO [78, 79] that were used previously in Ref. [57], using the same ultrasoft pseudopotentials that CASTEP generates on the fly for both codes. The valence electron configurations of $5s^{2}5p^{6}4f^{1}5d^{1}6s^{2}$ (with $4f$-electrons explicitly included) and $1s$ are adopted for the Ce and H atoms, respectively. Figure S11 in SI Sec. S7 shows close agreement between the three codes. Convergence tests at the DFT level yielded an energy cut-off of 800 eV for CASTEP and for Quantum ESPRESSO. For Wien2k, convergence is achieved when the product of the smallest atomic‐sphere radius and the maximum reciprocal‐space vector in the plane‐wave expansion reaches 7. A $\bm{\Gamma}$-centered $\mathbf{k}$-point grid of size $12\times 12\times 6$ is used for all three packages. The exchange-correlation functional is approximated using PBE [80]. As was done in previous works [12, 9, 47, 13, 57], we omit spin-orbit coupling for technical reasons. In the DFT + DMFT calculations, we use an on-site Coulomb interaction (Hubbard $U$) of $U=6.0$ eV and a Hund’s coupling $J=0.7$ eV for the cerium $f$ orbitals. This choice is comparable to previous values that were successfully used for elemental cerium [35, 36, 37, 38], CeH9 [57], and other Ce-compounds [44]. Note that we neglect any potential impact of pressure on Hubbard $U$ [41, 81]. We also omit any additional renormalizations of the DMFT electronic structure that arise from the electron-phonon coupling [82, 83].
Lattice dynamics calculations
@@
For the lattice dynamics calculations, we use finite difference with an in-house code which offers a versatile interface with DFT and DFT+DMFT [51], to calculate the phonon spectrum of CeH9. Considering that the computational cost of DFT+DMFT exhibits a particularly unfavorable scaling with the system size, we employ the non-diagonal supercell technique [84] to sample a $2\times 2\times 2$ $\mathbf{q}$-point grid in the Brillouin zone and then interpolate the result to obtain dynamical matrices on a $\mathbf{q}$-point grid of $6\times 6\times 3$. A convergence validation of the phonon dispersions with the size of the $\mathbf{q}$-point grid at the DFT level is provided in Sec. S8 of the SI. In addition, we find that the fixed self-energy approximation [51] is not applicable to CeH9, see Sec. S9 of the SI, and thus fully self-consistent DFT+DMFT calculations are performed for each displaced configuration. Specifically, we use four different supercells constructed with the following supercell matrices: $\big(\begin{smallmatrix}1&0&0\\ 0&1&0\\ 0&0&1\\ \end{smallmatrix}\big)$, $\big(\begin{smallmatrix}1&0&0\\ 0&1&0\\ 0&0&2\\ \end{smallmatrix}\big)$, $\big(\begin{smallmatrix}1&0&0\\ 0&2&0\\ 0&0&1\\ \end{smallmatrix}\big)$, and $\big(\begin{smallmatrix}1&0&0\\ 0&1&-1\\ 1&2&0\\ \end{smallmatrix}\big)$, the first one provides access to $\bm{\Gamma}-\mathbf{q}$-point, then the second one provides access to the $(0,0,1/2)$ $\mathbf{q}$-point, the third one to the $(0,1/2,0)$ $\mathbf{q}$-point, and the forth one to the $(1/2,1/2,1/2)$ $\mathbf{q}$-point. All other $\mathbf{q}$-points on the $2\times 2\times 2$ grid are related to these four by symmetry and therefore explicit calculations are not necessary.
Superconductivity calculations
@@
To evaluate the superconducting critical temperature $T_{\mathrm{c}}$, we used EPW [85, 86] to interpolate the el-ph coupling matrix calculated with the density functional perturbation theory as implemented in Quantum ESPRESSO from a $12\times 12\times 6$ $\mathbf{k}$-point grid and a $6\times 6\times 3$ $\mathbf{q}$-point grid to a $48\times 48\times 24$ $\mathbf{k}$-point grid and a $24\times 24\times 12$ $\mathbf{q}$-point grid. In particular, for $T_{\mathrm{c}}$ calculations at the DFT+DMFT level, the electronic and phononic degrees of freedom obtained from standard DFT are replaced with the counterparts obtained from DFT+DMFT. We used a standard value for the Morel-Anderson pseudopotential $\mu_{\mathrm{c}}=0.13$ which accounts for Coulomb repulsion of the system [58] in all our calculations.
@@
The data and codes that support the findings of this study are available in the Supplementary Information and on the Materials Cloud Archive at https://doi.org/10.24435/materialscloud:hw-nk.
@@
The Wien2k v21.1 code used in this study is a commercial software, available from http://susi.theochem.tuwien.ac.at. The Quantum ESPRESSO v7.0 code used in this research is open source: https://www.quantum-espresso.org. The CASTEP 23.1 code used in this study is freely available at https://www.castep.org for academic research. The EPW v5.4.1 code used in this research is open source: https://epw-code.org. The ana_cont v1.1.2 code used in this study is open source: https://github.com/josefkaufmann/ana_cont. Haule et al. provide their DMFT package (Oct-2024 release) at https://github.com/ru-ccmt/eDMFT.
@@
S.C. and Y.W. contributed equally to this work. S.C. performed the DMFT spectral function and superconducting calculations, while Y.W. performed the DMFT phonon calculations. S.P. and J.M.T provided technical help with the Wien2k, DMFT and EPW software. S.P., B.M. and J.M.T supervised the work. All authors reviewed and approved the final version of the manuscript.